# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 3.3.3. Summary
Document Visual Question Answering (Document VQA) is a relatively new field, pioneered by DocVQA (Mathew et al., 2021), which aims to generate or extract answers to natural language questions based on document images. This differs from key information extraction, which focuses on recognizing or extracting predefined key-value pairs. Document VQA requires a more comprehensive representation of the document and an understanding of correlations between the document and questions. Pretrained VRDU models (Xu et al., 2020b; Huang et al., 2022) demonstrate robust performance in single-page document understanding tasks. However, these models encounter challenges when applied to more typical and natural multi-page scenarios due to input length limitations. Recent solutions, such as (Tito et al., 2023; Blau et al., 2024), mainly address these challenges by identifying the page where a possible answer may be located and then applying SoTA single-page techniques to retrieve the answer. Despite these advancements, real-world applications often present more complex situations, such as long-term dependencies and cross-page entity relationships, which still need further exploration in the document VQA field.
4. Multi-Task VRD Understanding Models
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Multi-task based models, for tree=fill=brown!45 [Fine-grained Pretrained Models (Encoder-only), for tree=fill=violet!40 [Layout-aware Pretrained Models, for tree=fill=violet!20 [LayoutLM (Xu et al., 2020a); BROS(Hong et al., 2021); LiLT (Wang et al., 2022b); XDoc (Chen et al., 2022); LayoutMask (Tu et al., 2023); StructuralLM (Xu et al., 2020a), for tree=fill=violet!10] ] [Visual-integrated Pretrained Models, for tree=fill=violet!20 [LayoutLMv2(Xu et al., 2020b); LayoutXLM(Xu et al., 2021); DocFormer(Appalaraju et al., 2021); LayoutLMv3(Huang et al., 2022), for tree=fill=violet!10] ] ] [Coarse and Joint-grained Pretrained Models (Encoder-only), for tree=fill=pink!40 [Coarse-grained Models, for tree=fill=pink!20 [SelfDoc(Li et al., 2021b); UniDoc(Gu et al., 2021)], for tree=fill=pink!10 ] [Joint-grained Models, for tree=fill=pink!20 [StrucText(Li et al., 2021c); Fast-StrucText(Zhai et al., 2023); MGDoc (Wang et al., 2022a); WUKONG-READER (Bai et al., 2022); GeoLayoutLM (Luo et al., 2023), for tree=fill=pink!10] ] ] [Encoder-Decoder Pretrained Frameworks, for tree=fill=red!40 [OCR-dependent Encoder-Decoder Frameworks, for tree=fill=red!20 [TiLT(Li et al., 2021b); UDOP(Tang et al., 2022); DocFormerv2(Appalaraju et al., 2024); ViTLP(Mao et al., 2024), for tree=fill=red!10] ] [OCR-independent Frameworks, for tree=fill=red!20 [Donut(Kim et al., 2022); Dessurt(Davis et al., 2022); ReRum (Cao et al., 2023a); StructTextV2 (Yu et al., 2022), for tree=fill=red!10] ] [LLM-based Frameworks, for tree=fill=red!20 [HRVDA(Liu et al., 2024b); LayoutLLM(Luo et al., 2024), for tree=fill=red!10] ] ] [Non-Pretrained Frameworks, for tree=fill=yellow!40 [CALM(Du et al., 2022); LayoutGCN(Shi et al., 2023),for tree=fill=yellow!20 ] ] ]
@@
Figure 3. Multi-task visually rich document understanding frameworks.
Models designed for specific VRDU tasks often incorporate task-oriented techniques. Drawing inspiration from pretrained models in the vision (Dosovitskiy et al., 2020; Ren et al., 2015) and language (Devlin, 2018; Liu et al., 2019b) domains, enhancing document representations may significantly improve the performance of various downstream tasks. Consequently, a range of models have been developed to extract knowledge from extensive document collections. These models employ different pretrained tasks based on their architecture, the modalities they process, and the granularity of the information they extract from documents. Moreover, some models introduce specialised techniques to boost the effectiveness of pretrained model representations, achieving better performance without heavy pretraining. This section explores various document understanding frameworks focused on enhancing document representation for robustness and comprehensiveness in addressing multiple VRDU downstream tasks.
4.1. Fine-grained Pretrained Models
@@
4.1.1. Layout-aware Pretrained Language Modeling
@@
LayoutLM (Xu et al., 2020a) is the first pretrained document understanding model by leveraging textual and layout information in the pretraining stage. BERT architecture is the backbone and 2-D positional embedding  [^3] Please refer to Section xx for more detailed information on 2-D positional encoding with textual information is used to pretrain on IIT-CDIP Test Collection 1.0. Two specific pretraining tasks are first introduced, named Masked Visual-Language Model (MVLM) and Multi-label Document Classification, to generate layout-aware textual representation and more comprehensive document representation, respectively. Like Masked Language Modeling adopted by most pretrained language models, MVLM randomly masks some input tokens but keeps the corresponding 2-D position embeddings to predict masked tokens to ensure the pretrained model is aware of the spatial relations between input tokens. MDC is a supervised pretraining task to predict the input document types (e.g. forms, exam papers, academic papers) that generate more comprehensive document-level representations. The fine-tuned LayoutLM could perform much better than textual-only frameworks on key information extraction (Jaume et al., 2019; Huang et al., 2019) and document classification (Harley et al., 2015).
BROS (Hong et al., 2021) aims to propose a pretrained VRDU model to represent the continuous property of 2D space by introducing a new 2-D positional encoding and a textual-spatial correlation aware attention score to replace the vanilla self-attention. Supposing $[x1,y1,x2,y2,x3,y3,x4,y4]$ is the bounding box coordinates of input token $t$, $p^{1}=[\mathcal{F}_{sin}(x1)\oplus\mathcal{F}_{sin}(y1)],p^{1}\rightarrow\mathbb{R}^{d_{pos}}$. The final positional representation of $t$ is $pos_{t}=W_{p1}p_{1}+W_{p2}p_{2}+W_{p3}p_{3}+W_{p4}p_{4}$, where all $W_{p}\in\mathbb{R}^{2d_{pos}\times d}$. The new 2-D positional encoding intends to provide a more natural way to encode the continuous 2-D coordinates. Additionally, instead of simply summing textual and positional representations, a novel attention score is introduced to consider both textual and spatial features and their correlations. To calculate the attention score $\alpha_{ij}$ between $t_{i}$ and $t_{j}$, they consider correlation intra and inter-correlation between textual and positional modalities.
(12) $\alpha_{ij}=(W^{q_{t_{i}}}T_{i})^{\top}(W^{q_{t_{j}}}T_{j})+(W^{q_{t_{i}}}T_{i}\circ W^{pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{pos_{t_{j}}}pos_{t_{j}})+(W^{\prime pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{\prime pos_{t_{j}}}pos_{t_{j}})$
where $(W^{q_{t_{i}}}T_{i})^{\top}(W^{q_{t_{j}}}T_{j})$ and $(W^{\prime pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{\prime pos_{t_{j}}}pos_{t_{j}})$ is the intra-modality attention scores of textual and positional modalities between two tokens. $(W^{q_{t_{i}}}T_{i}\circ W^{pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{pos_{t_{j}}}pos_{t_{j}})$ is used to formulate the spatial dependency given the source semantic representation. Additionally, inspired by SpanBERT (Joshi et al., 2020), they use an area-masked language model to mask tokens by random-sized rectangle regions.
StructuralLM (Li et al., 2021a) is the first VRUD model using image patches (named ”cells” in the paper) to group input tokens for conducting various pretraining tasks. They use BERT as the backbone and generate multimodal representations $P$ of each patch, which will be used for two pretraining tasks, MVLM and Cell (Patch) Position Classification (CPC). The bounding box $(x0,y0,x1,y1)$ of an image patch is encoded by the 2-D positional encoding introduced by LayoutLM and the $n$ tokens inside $p$ are represented as $\{t_{1},t_{2},\dots,t_{n}\}$ which share the same 2-D positional encoding as for tokens belongs to one patch. A token $t_{i}$ could be represented as summing up token representation $T_{i}$, 2-D positional encoding $pos_{t_{i}}^{2d}$ and 1-D position embedding $pos_{i}$. Two pretraining tasks are used to understand the patch-level spatial correlations, including MVLM and CPC. MVLM is similar to LayoutLM but uses patch-level layout embeddings instead of token-level. Another novel cell position classification task is applied to predict the area index of randomly selected tokens from $N$ evenly split areas. Two pretraining tasks are performed simultaneously to capture the patch-level spatial dependencies between input tokens.
LiLT (Wang et al., 2022b) introduces a language-independent layout Transformer for mono/multi-lingual document understanding. The text and layout information are first separately encoded and jointly learned during pretraining. In the fine-tuning stage, two modality representations are concatenated to perform downstream tasks. The input token representations follow BERT, where $j$-th token is $T_{j}=T_{j}+pos_{t_{j}}+pos_{t_{j}}^{2d}$. The layout representations are slightly different from LayoutLM, normalising the bbox coordinates in the $[0,1000]$ and four linear layers encode the x-axis, y-axis, width and height features. The normalised bbox of a token $t$ is $[x0,y0,x1,y1,w,h]$ is encoded to get the final layout representation $L$:
(13) $L=W_{L}(W_{x}x0\oplus W_{y}x0\oplus W_{x}x1\oplus W_{y}y1\oplus W_{w}w\oplus W_{h}h)+pos_{L}$
where $W_{L}\in\mathbb{R}^{6d_{L}^{\prime}\times d_{L}}$ and $W_{x},W_{y},W_{w},W_{h}\in\mathbb{R}^{1\times d_{L}}$. The text and layout embedding are fed into two sub-models to generate high-level representations. A bi-directional attention complementation mechanism (BiACM) is proposed to augment the cross-modality interaction  [^4] We provide a detailed explanation of BiACM in Section X. Three self-supervised learning methods are proposed to enable the model to understand the multimodal document representation, including MVLM and Key Point Location, which is similar to the CPC proposed by StructureLM to predict the area index of masked tokens based on the layout features. Additionally, a text and layout alignment task is proposed to enhance further the cross-modality representation to predict whether each pair is aligned. LiLT could achieve promising performance on multi-lingual document understanding benchmarks (Wang et al., 2021a; Xu et al., 2021).
XDoc (Chen et al., 2022) propose a unified framework to deal with text inputs from multi-format inputs, including plain text, document and web text. Various encoding methods are proposed to tackle diverse text formats. For plain text token $t_{pln}$, the token representation follows BERT to get $T_{pln}$. For document text token $t_{doc}$, they adopted similar strategies (Xu et al., 2020a; Hong et al., 2021) to get $T_{doc}$ by summing up initial token representation $E_{t_{doc}}$, 1-D position embedding $pos_{t_{doc}}$ and 2-D position embedding $pos_{t_{doc}}^{2d}$. $pos_{t_{doc}}^{2d}$ uses different bbox formats $[x_{0},x_{1},y_{0},y_{1},w,h]$ following a Linear-ReLu-Linear adaptor to make the document format text embedding more representative. To encode web text inputs, they follow the way introduced by MarkupLM to encode the source file tags and subscript information to get the XPath embedding. The same structured adaptor is leveraged for better pretraining. All three text formats are pretrained by MLM and fine-tuned on plain text (Rajpurkar et al., 2016a; Wang et al., 2018), document (Jaume et al., 2019; Mathew et al., 2021), and web text (Chen et al., 2021) benchmark datasets, respectively.
LayoutMask (Tu et al., 2023) aims to improve the text-layout interactions by using local 1D positional encoding and new pretraining tasks. OCR-dependent document understanding frameworks are suffered from improper reading orders. It uses LayoutLMv2 as the backbone but removes the visual module, and the local 1D positional encoding is adopted to replace the global 1D positional encoding, which only encodes the token ordered within each segment detected by OCR tools and always restarts with 1 from each individual segment. Moreover, the 2D positional encoding is also using segment bbox instead of individual words. The first task pertains to traditional masked language modelling, but two different masking strategies are used. Firstly, they set the mask at the word level instead of masking the token itself to enable the model to capture more contextual information to predict the masked word. Additionally, to boost cross-segment understanding, the probability of masking the first and last word of a segment is higher than that of others. Another Masked Position Modelling is proposed to ask the model to predict the masked word-level 2D positions to promote the layout information representation.
4.1.2. Visual Integrated Models
Integrating visual cues alongside text and layout information during the pretraining stage can significantly enhance model capabilities, capturing more comprehensive document insights than text and layout alone. Various frameworks and pretraining tasks focusing solely on text and layout have been expanded to include visual-text matching tasks to strengthen cross-modality alignment. This integration enables models to interpret better the complex interplay of visual, textual, and layout features within documents. Existing models that utilise these comprehensive inputs can be categorised based on their approach to visual feature acquisition: feature map-based models, which generate comprehensive visual representations, and patch pixel-based models, which focus on granular visual details. This classification helps understand how different models leverage visual information to enhance document understanding.
LayoutLMv2 (Xu et al., 2020b) is the first pretrained model integrating text, layout and visual aspects during the pretraining stage. A single framework multimodal transformer is applied to get the visual, text and layout features simultaneously, where each input textual and visual tokens are assigned a segment ID to distinguish the modality or semantic type. The textual and layout representations generally follow LayoutLM, except the linear projected segment ID, $seg_{t}$, embeddings are addicted to original textual embeddings. The input document image is firstly fed into a trainable visual encoder (based on ResNeXt-FPN), of which output is evenly split into $m$ flattened visual tokens, $v_{token}$ following a linear layer to project into the same dimension as textual embedding. The $j$-th visual token embedding is $V_{j}=W_{v_{j}}+pos_{j}+W_{seg_{v}}$. Moreover, to capture the relative position information of inter/intra-modality features, spatial-aware self-attention is introduced by adding bias terms of 1-D ($b^{1D}$) and 2-D ($b^{2D}$) relative position. The 1-D relative positional bias between input visual or textual tokens $t_{i}$ and $t_{j}$ is $b^{1D}=W_{b_{1D}}(j-i)$ and $b^{2D}=b^{2D}_{x}+b^{2D}_{y}$, where $b^{2D}_{x}=W_{b^{2D}_{x}}(x0_{i}-x0_{j})$, $b^{2D}_{y}=W_{b^{2D}_{y}}(y0_{i}-y0_{j})$. LayoutLMv2 is pretrained on MVLM and the other two new proposed tasks, Text-Image Alignment (TIA) and Text-Image Matching (TIM)  [^5] Please refer to Section to check more detailed information about those two pretraining tasks. to capture more cross-modality alignment.
LayoutXLM (Xu et al., 2021) extends the LayoutLMv2 architecture to a multilingual setup where the MVLM is extended to Multilingual Mased Visual-Language Modeling pretrained on 22 million self-collected digital-born PDF files and 8 million scanned English documents from IIT-CDIP processed by off-the-shelf PDF parsers.
DocFormer (Appalaraju et al., 2021) is a multimodal transformer encoder-based architecture which also uses a trainable CNN-backbone (ResNet50) to extract the visual cues of input document image to get visual representations $\mathbb{V}\in\mathbb{R}^{(d\times N)}$, where $d=768$ is the transformer hidden state and $N=512$ is the number of visual tokens. The initial textual representations $\mathbb{T}$ are acquired by feeding the OCR extracted text with bbox into LayoutLM where $\mathbb{T}\in\mathbb{R}^{(d\times N)}$. 2D positional encoding is also used by using $W^{x}_{v},W^{y}_{v}$ and $W^{x}_{t},W^{y}_{t}$ to encode $x,y$ axis, as well as, visual and textual aspects to acquire $pos^{2d}_{v}$ and $pos^{2d}_{t}$. More positional features are adopted such as width, height and relative distance between neighbours. Moreover, certain inductive biases are applied to get a new self-attention score $\alpha$ for paying more attention to local features. Supposing $\alpha_{ij}^{v}$ is attention score between visual feature $V_{i}$ and $V_{j}$, we could have $\alpha_{ij}^{v}=(W^{K}_{v}V_{j})^{\top}(W^{Q}_{v}V_{i})+(pos_{ij})^{\top}W^{Q}_{v}V_{i}+(pos_{ij})^{\top}W^{K}_{v}V_{j}+(W^{Q}_{s}pos_{ij}^{2d})^{\top}(W^{K}_{s}pos_{ij}^{2d})$, where $pos_{ij}$ is the 1D relative positional encoding. The same procedures are applied to textual features as well, and two modalities share the same $W^{Q}_{s}$ and $W^{K}_{s}$ to help the model correlate features across modalities. The multi-modal token representations $\mathbb{M}$ of $l$ encoder layer is $\mathbb{M}_{l}=\mathbb{T}_{l}+\mathbb{V}_{l}$.
LayoutLMv3 (Huang et al., 2022) is the first pretrained VRDU model to encode visual features without using heavy CNN-backbones. LayoutLMv3 uses the exact same way as LayoutLMv2 to encode the textual and layout information except replacing the word-level bbox to segment-level for 2-D positional embedding. For visual feature representation, they follow ViT and ViLT to linearly project the evenly split image patches with learnable 1-D positional encoding to the transformer encoder. For the pretraining setting, they use masked language modelling to mask 30% tokens drawn from a Poisson distribution. To augment the visual representation by contextually learning with multimodal features, a Masked Image Modelling (MIM) (Bao et al., 2021) task is used to mask 40% image tokens with clockwise masking randomly. Each image token will be converted into discrete tokens following (Ramesh et al., 2021), and a cross-entropy loss is adopted to reconstruct the masked image tokens. Moreover, to explicitly learn the correlation between text and image modalities, a Word-Patch Alignment (WPA) pretraining objective is applied. WPA is a binary classification task to predict whether the unmasked token-patch pair is ”aligned” or ”unaligned”.
4.2. Coarse and Joint-grained Pretrained Models

### v2

@@ 3.3.3 Summary of VRD-QA Models
VRD-QA is a relatively new research area, introduced by DocVQA [docvqa], which focuses on answering natural language questions based on document images. Unlike key information extraction—which targets predefined key-value pairs—VRD-QA requires a deeper understanding of the entire document and its relevance to the question, demanding more comprehensive semantic and structural representations. Pretrained VRD-CU models such as LayoutLMv2 and LayoutLMv3 [layoutlmv2, layoutlmv3] perform well on single-page document QA tasks. However, they face limitations when applied to multi-page documents due to input length constraints. To address this, recent works [mpdocvqa, gram] have proposed identifying the most relevant page(s) and then applying single-page QA techniques to extract answers. Despite these advances, real-world applications often involve more complex scenarios—such as long-range dependencies and cross-page relationships—which remain open challenges in the VRD-QA field and warrant further investigation.
3.4 Summary of Mono-Task Models
To achieve better performance on specific VRD-CU tasks, mono-task frameworks often adopt task-specific designs—for example, graph-based spatial encoding for KIE [formnet, formnetv2] or page-locating modules for multi-page VRD-QA [mpdocvqa]. While optimized for individual objectives, these specialized components often form the foundation for multi-task learning. For instance, relation-aware modeling in KIE can support entity linking, and layout-aware encoders used in QA can be adapted for structured extraction. Recognizing these shared components is essential for understanding cross-task generalization and designing unified VRD-CU systems. Early frameworks relied heavily on hand-crafted features and rule-based heuristics, such as bounding box grouping or document-specific layout parsing. However, with the emergence of pretrained models like LayoutLM, BROS, and LiLT [layoutlm, bros, lilt], these manual designs have been largely replaced by shared multimodal encoders trained on large-scale document corpora. This shift has enabled the development of more general, instruction-tuned frameworks capable of handling multiple tasks with minimal adaptation. Section 3 builds on this trend, introducing multi-task models that unify these components through shared architectures, input granularity settings, and cross-task generalization strategies.
4 Multi-Task VRD Understanding Models
@@
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Multi-task based models, for tree=fill=brown!45 [Fine-grained Pretrained Models (Encoder-only), for tree=fill=violet!40 [Layout-aware Pretrained Models, for tree=fill=violet!20 [LayoutLM [layoutlm]; BROS[bros]; LiLT [lilt]; XDoc [xdoc]; LayoutMask [layoutmask]; StructuralLM [layoutlm], for tree=fill=violet!10] ] [Visual-integrated Pretrained Models, for tree=fill=violet!20 [LayoutLMv2[layoutlmv2]; LayoutXLM[layoutxlm]; DocFormer[docformer]; LayoutLMv3[layoutlmv3], for tree=fill=violet!10] ] ] [Coarse and Joint-grained Pretrained Models (Encoder-only), for tree=fill=pink!40 [Coarse-grained Models, for tree=fill=pink!20 [SelfDoc[selfdoc]; UniDoc[udoc]], for tree=fill=pink!10 ] [Joint-grained Models, for tree=fill=pink!20 [StrucText[structext]; Fast-StrucText[faststructext]; MGDoc [mgdoc]; WUKONG-READER [wukong]; GeoLayoutLM [geolayoutlm], for tree=fill=pink!10] ] ] [Encoder-Decoder Pretrained Frameworks, for tree=fill=red!40 [OCR-dependent Encoder-Decoder Frameworks, for tree=fill=red!20 [TiLT[selfdoc]; UDOP[udop]; DocFormerv2[docformerv2]; ViTLP[vitlp], for tree=fill=red!10] ] [OCR-free Frameworks, for tree=fill=red!20 [Donut[donut]; Dessurt[dessurt]; SeRum [serum]; StructTextV2 [structextv2], for tree=fill=red!10] ] [LLM-based Frameworks, for tree=fill=red!20 [HRVDA[hrvda]; LayoutLLM[layoutllm], for tree=fill=red!10] ] ] [Non-Pretrained Frameworks, for tree=fill=yellow!40 [CALM[calm]; LayoutGCN[layoutgcn],for tree=fill=yellow!20 ] ] ]
@@
Figure 3: Multi-task visually rich document understanding frameworks.
As illustrated in Figure 3, recent frameworks for VRD-CU increasingly adopt multi-task learning designs. These frameworks can be broadly categorized based on their model architecture and input granularity, which are key to determining their capability and generalizability. In terms of architecture, models are typically classified as Encoder-only, Encoder-Decoder. Encoder-only models utilize transformer-based encoders to capture multimodal document features and are often self-supervised pretrained on large-scale datasets. Encoder-Decoder models, usually generation-oriented, consist of an encoder that processes visual or textual input and a decoder that produces structured output. These can be further divided into OCR-dependent models, which require external text input, and OCR-free models that operate directly on document images. Non-pretrained models, in contrast, are task-specific and avoid extensive pretraining, often using lightweight or graph-based architectures. Input granularity distinguishes how models represent document information. Fine-grained models encode detailed features at the token or region level, preserving spatial and textual fidelity. Coarse-grained models abstract over larger structural units, such as blocks or paragraphs, to capture high-level semantics. Joint-grained approaches integrate both levels to support comprehensive document understanding.
4.1 Fine-grained Pretrained Models
@@
4.1.1 Layout-aware Pretrained Language Modeling
@@
LayoutLM [layoutlm] is the first pretrained document understanding model by leveraging textual and layout information in the pretraining stage. BERT architecture is the backbone and 2-D positional embedding  [^1] Please refer to Section 6.1.3 for more detailed information on 2-D positional encoding with textual information is used to pretrain on IIT-CDIP Test Collection 1.0. Two specific pretraining tasks are first introduced, named Masked Visual-Language Model (MVLM) and Multi-label Document Classification, to generate layout-aware textual representation and more comprehensive document representation, respectively. Like Masked Language Modeling adopted by most pretrained language models, MVLM randomly masks some input tokens but keeps the corresponding 2-D position embeddings to predict masked tokens to ensure the pretrained model is aware of the spatial relations between input tokens. MDC is a supervised pretraining task to predict the input document types (e.g. forms, exam papers, academic papers) that generate more comprehensive document-level representations. The fine-tuned LayoutLM could perform much better than textual-only frameworks on key information extraction [funsd, sroie] and document classification [rvlcdip].
BROS [bros] proposes a pretrained VRDU model that captures the continuous nature of 2D space through a novel positional encoding and a correlation-aware attention mechanism. Given a token’s bounding box coordinates $[x_{1},y_{1},\dots,x_{4},y_{4}]$, the positional embedding is computed as:
$\text{pos}_{t}=W_{p1}p_{1}+W_{p2}p_{2}+W_{p3}p_{3}+W_{p4}p_{4},$
where each $p_{i}=\mathcal{F}_{\sin}(x_{i})\oplus\mathcal{F}_{\sin}(y_{i})$ and $W_{p*}\in\mathbb{R}^{2d_{\text{pos}}\times d}$.
To replace vanilla self-attention, BROS introduces an attention score $\alpha_{ij}$ that captures intra- and inter-modal interactions between textual and positional features:
$\alpha_{ij}=$ $\underbrace{\left(W^{q}T_{i}\right)^{\top}\left(W^{q}T_{j}\right)}_{\text{text-text}}+\underbrace{\left(W^{q}T_{i}\circ W^{\text{pos}}\text{pos}_{i}\right)^{\top}\left(W^{\text{pos}}\text{pos}_{j}\right)}_{\text{text-pos}}$
$+\underbrace{\left(W^{\prime\text{pos}}\text{pos}_{i}\right)^{\top}\left(W^{\prime\text{pos}}\text{pos}_{j}\right)}_{\text{pos-pos}}.$ (9)
Inspired by SpanBERT [joshi2020spanbert], BROS also adopts an area-masked language modeling task, masking tokens within randomly sampled rectangular regions to enhance span-level understanding.
StructuralLM [structurallm] is the first VRDU model to leverage image patches (termed ”cells”) to group tokens and perform patch-level pretraining. Built on BERT, it computes multimodal patch representations $P$ and introduces two pretraining tasks: Masked Visual-Language Modeling (MVLM) and Cell Position Classification (CPC). Each patch is defined by a bounding box $(x_{0},y_{0},x_{1},y_{1})$ and encoded using the 2-D positional encoding from LayoutLM. Tokens $\{t_{1},t_{2},\dots,t_{n}\}$ within a patch share this 2-D position. A token $t_{i}$ is represented by: $t_{i}=T_{i}+\text{pos}^{2D}_{t_{i}}+\text{pos}^{1D}_{i},$ where $T_{i}$ is the token embedding, $\text{pos}^{2D}_{t_{i}}$ is the patch-level layout embedding, and $\text{pos}^{1D}_{i}$ is the token’s sequence position. MVLM masks tokens using patch-level layout embeddings, while CPC predicts the area index of a token among $N$ equally divided regions. Together, these tasks encourage the model to capture spatial dependencies at the patch level.
LiLT [lilt] proposes a language-independent layout Transformer for mono- and multilingual document understanding. Text and layout are encoded separately and fused during pretraining, with their representations concatenated for downstream tasks. Each token $t_{j}$ is represented as:
$T_{j}=T_{j}+\text{pos}_{t_{j}}^{1D}+\text{pos}_{t_{j}}^{2D}.$
Unlike LayoutLM, layout features are normalised to $[0,1000]$, and six layout attributes $[x_{0},y_{0},x_{1},y_{1},w,h]$ are encoded via:
$L=W_{L}(W_{x}x_{0}\oplus W_{y}y_{0}\oplus W_{x}x_{1}\oplus W_{y}y_{1}\oplus W_{w}w\oplus W_{h}h)+\text{pos}_{L},$
where $W_{L}\in\mathbb{R}^{6d^{\prime}_{L}\times d_{L}}$ and $W_{*}$ are learnable projections. Text and layout embeddings are processed via two sub-models, with a bi-directional attention complementation mechanism (BiACM) enhancing cross-modal interaction. Three pretraining tasks are introduced: MVLM, Key Point Location (predicting token area indices), and a text-layout alignment task (predicting whether text-layout pairs match). LiLT achieves strong results on multilingual document benchmarks [ephoie, layoutxlm].
XDoc [xdoc] presents a unified architecture that handles diverse input formats—plain text, documents, and web data. It applies different encoding strategies: BERT-style for plain text; LayoutLM-style for documents with adapted 2D box encoding; and XPath-based embedding for web structures. All formats are pretrained with masked language modeling (MLM) and fine-tuned on corresponding benchmarks.
LayoutMask [layoutmask] targets the issue of improper reading order in OCR-dependent systems. It removes visual inputs and replaces global 1D positional encoding with a segment-based local variant, restarting the order within each OCR segment. The model introduces two pretraining objectives: (1) word-level MVLM with higher masking probabilities for segment boundaries, and (2) Masked Position Modeling, predicting the 2D coordinates of masked words to reinforce layout understanding.
4.1.2 Visual Integrated Models
Integrating visual cues with textual and layout information during pretraining significantly enhances a model’s ability to capture rich and comprehensive document semantics. While earlier frameworks primarily focused on text and layout, recent approaches extend these by introducing visual-text alignment tasks to strengthen cross-modal understanding. This integration allows models to better interpret the intricate relationships among visual elements, textual content, and spatial structures within documents. Existing vision-integrated models can be broadly categorized based on how visual features are extracted: (1) RoI-based models, which employ region-of-interest (RoI) alignment using CNN backbones to generate high-level visual representations; and (2) pixel-based models, which directly extract pixel-level features via Vision Transformers.
LayoutLMv2 [layoutlmv2] is the first pretrained model to integrate textual, layout, and visual modalities into a unified multimodal Transformer. Building upon LayoutLM, it introduces a trainable ResNeXt-FPN visual encoder whose outputs are projected to the same dimensional space as textual embeddings. Each modality is assigned a segment ID (e.g., $seg_{t}$, $seg_{v}$) added to its embeddings. Spatial-aware self-attention is adopted, incorporating 1D and 2D relative positional biases:
$b^{1D}$ $=W_{b_{1D}}(j-i),$ (10)
$b^{2D}$ $=W_{b^{2D}_{x}}(x0_{i}-x0_{j})+W_{b^{2D}_{y}}(y0_{i}-y0_{j})$ (11)
Three pretraining tasks are used: MVLM, Text-Image Alignment (TIA), and Text-Image Matching (TIM), enhancing cross-modal understanding.
LayoutXLM [layoutxlm] extends LayoutLMv2 to multilingual settings with the same architecture but trains on 22M multilingual PDF and 8M scanned English documents. The MVLM task is adapted for multilingual contexts.
DocFormer [docformer] proposes a multimodal encoder architecture combining LayoutLM-based text encoding and CNN-based visual encoding via ResNet-50. Textual ($\mathbb{T}\in\mathbb{R}^{d\times N}$) and visual ($\mathbb{V}\in\mathbb{R}^{d\times N}$) features are enhanced with 2D and 1D positional embeddings. A novel self-attention score $\alpha_{ij}^{v}$ includes multiple spatial and semantic biases:
$\alpha_{ij}^{v}$ $=(W^{K}_{v}V_{j})^{\top}(W^{Q}_{v}V_{i})+(pos_{ij})^{\top}W^{Q}_{v}V_{i}+(pos_{ij})^{\top}W^{K}_{v}V_{j}$
$\quad+(W^{Q}_{s}pos_{ij}^{2d})^{\top}(W^{K}_{s}pos_{ij}^{2d})$ (12)
Two modalities are fused at each encoder layer via addition: $\mathbb{M}_{l}=\mathbb{T}_{l}+\mathbb{V}_{l}$.
LayoutLMv3[layoutlmv3] enhances model efficiency by eliminating heavy CNNs and adopting a pure Transformer-based vision encoder. Inspired by ViT, it divides document images into patches and projects them linearly with learnable 1D positional encodings. Layout information is encoded using segment-level bounding boxes with 2D positional embeddings. In addition to masked language modeling (MLM), LayoutLMv3 introduces two new pretraining tasks: Masked Image Modeling (MIM), where 40% of image patches are masked and reconstructed using discrete tokens with a cross-entropy loss[bao2021beit]; and Word-Patch Alignment (WPA), a binary classification objective that predicts whether a text token and image patch pair is aligned. Together, these tasks promote fine-grained multimodal alignment between visual and textual features.
4.2 Coarse and Joint-grained Pretrained Models